Derive area of parallelogram from the pair of homogenous and nonhomogeneous straight lines

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
Lets Consider parallelogram ABCD where height =h
after joining BD diagonal, we find triangle ADB and CDB to be congruent (SSS)
so area of CDB and ADB are same
area of parallelogram ABCD =area of CDB+area of ABD
so ar of parallelogram =2*ar triangle CDB
ar triangle CDB= 1/2*base*ht=1/2*CD*h
therefore ar of parallelogram =2* 1/2* CD* h =CD*h
or area of parallelogram =base * ht     [ans]
          
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)